OpenStudy (anonymous):

The sum of three consecutive even integers is 108. What is the largest number? 34 35 37 38

OpenStudy (anonymous):

-34, -36, -38 108/3 = 36 36+2 = 38 36-2 = 34

OpenStudy (anonymous):

Good job citizen. @brittany210

OpenStudy (anonymous):

Thank you @LearningIsAwesome

OpenStudy (abmon98):

first term n second term n+1 third term n+2 n+(n+1)+(n+2) 3n+3=108 105=3n 105/3=n n=35

OpenStudy (anonymous):

thank you so its 34

OpenStudy (abmon98):

35+36+37=108 so the largest number is 37

OpenStudy (mathmate):

@keoniawhite It always pays to reread the question before writing the final answer!

OpenStudy (mathmate):

@Abmon98 It always pays to reread the question before writing the final answer! (although I miss sometimes too!)

OpenStudy (abmon98):

oh even** i am sorry yes you are right i thought its 3 consecutive integers.

OpenStudy (anonymous):

so i am still lost can one of yall help me get to the answer

OpenStudy (mathmate):

Can you post the question (up to the question mark) again but do not do copy and paste?

OpenStudy (mathmate):

@keoniawhite

OpenStudy (anonymous):

the sum of three consecutive even integers is 108. What is the largest number?

OpenStudy (anonymous):

do you need the answer choices?

OpenStudy (mathmate):

In mathematics questions, every word counts.

OpenStudy (anonymous):

thats the whole ? i wrote it again

OpenStudy (mathmate):

Yes, but you have to take into account of every word of the question. What was your proposed answer again? @keoniawhite

OpenStudy (anonymous):

i thought it was 37 but i really dont know the answer

OpenStudy (mathmate):

Was 37 the largest number of the three? Was 37 among the three consecutive even integers?

OpenStudy (anonymous):

i figured it out thank you for yall help

OpenStudy (mathmate):

You're welcome! :)